Given the following annual returns for Stock ABC, what does a $1,500 investment grow to over the 5-year period?
Year 1: 15%
Year 2: 20%
Year 3: 25%
Year 4: 15%
Year 5: 15%
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period

A=1500*(1.15)*(1.2)*(1.25)*(1.15)*(1.15)

=1500*2.2813125

=$3421.97(Approx).